# CLUENet: Cluster Attention Makes Neural Networks Have Eyes

### v1

@@ Introduction
Results on CIFAR-100. As shown in Table 1, in the convolution paradigm, ShuffleNet_v2 achieves the best Top-1 accuracy of 71.09%. For the attention paradigm, PVT_v2 leads with 70.77%. Among clustering-paradigm, CLUENet stands out with a Top-1 accuracy of 76.55%, surpassing existing clustering model CoC by 4.63% and outperforming PVT_v2 and ShuffleNet by 5.78% and 5.46%, respectively.
Results on Mini-ImageNet. Table 2 summarizes the classification results on the Mini-ImageNet validation set. Under similar parameter scales, CLUENet (base) achieves the best performance among clustering-paradigm models, surpassing CoC (medium), FEC (large), and ClusterFormer (tiny) by 4.05%, 3.11%, and 8.45% in Top-1 accuracy, and by 2.18%, 1.92%, and 4.21% in Top-3 accuracy, respectively. Without using windowing mechanisms while adopting a relatively dense cluster count, CLUENet (base) achieves an inference speed of 679.05 img/s, which is comparable that of FEC (550.69 img/s) and CoC (605.67 img/s), demonstrating its efficient architecture and strong feature representation.
Comparing with the Attention-based models, CLUENet (micro) outperforms the attention-based PVT_v2 (b0) by 3.41% in Top-1 accuracy and 2.31% in Top-3 accuracy while using 0.42M fewer parameters. It also surpasses EfficientFormer (s0) by 0.83% in Top-1 and 1.12% in Top-3 accuracy with a similar parameter budget. CLUENet (tiny) outperforms EfficientFormer (s1) by 1.54% in Top-1 and 0.97% in Top-3 accuracy. Furthermore, CLUENet (small) surpasses PVT_v2 (b1) and EfficientFormer (s1) by 3.92% and 1.74% in Top-1, respectively.
Compared with the convolution-based models, CLUENet (micro) achieves slightly higher Top-1 and Top-3 accuracy than ShuffleNet_v2 (x1.5) (+0.36%, +0.43%), and CLUENet (tiny) outperforms ShuffleNet_v2 (x2.0) and ResNet18 by 0.88% and 3.56% in Top-1 accuracy, and by 0.36% and 1.41% in Top-3 accuracy, respectively. Moreover, CLUENet (small) surpasses both ResNet18 and ConvNeXt_v2 (N) by 4.54% and 6.45% in Top-1 accuracy, respectively.
@@
Model Visualization. The proposed CLUENet offers intrinsic interpretability through its architecture. We visualize the cluster receptive fields of CLUENet (tiny) across all four stages on the MiNi-ImageNet classification task. As shown in Fig. 5, the model captures multi-level semantic structures at different stages, reflecting a hierarchical feature extraction process similar to human visual perception. For clarity, only one semantic head is visualized per stage. Additionally, we visualize the receptive field of each position on the $7\times 7$ feature map from all fourth stages to highlight the semantic areas most influential for classification. Since directly displaying all 49 clusters can be overwhelming, we adopt the $K$-Means-based cluster merging method proposed by Chen et al. (ref19) to reduce the number of visualized clusters. To preserve the interpretability purity of the network, we visualize the fourth-stage $7\times 7$ feature map with increasing cluster numbers by adjusting the K-Means merge parameter using the K-Means.
@@
Figure 6: Visualization results of different paradigm models. ResNet shows Grad-CAM activation maps at four stages. PVTv2 presents attention maps based on the center key token in the first attention head of the last block at each stage. CoC, ClusterFormer and FEC visualize one attention head from one block per stage. FEC and CLUENet additionally provide the final-stage semantic clustering maps used for classification.
As illustrated in Fig. 5 (a), CLUENet exhibits a progressive semantic understanding. Early stages (1–3) focus on the object’s overall contour and location, separating it from the background to form clear object boundaries. The deeper stages (3–4) progressively shift attention to finer and more discriminative regions critical for classification. For example, in the first row of Fig. 5, the model initially treats the bird and branch as a whole, but by Stage 3, it distinguishes the branch as background and focuses on the bird’s head and torso in Stage 4. As shown in Fig. 5 (b), when the number of clusters is further increased, the same semantic object is split into multiple semantic parts, capturing finer details of the object. Importantly, all visualized samples are correctly classified, and the final-stage feature maps (the last column of Fig. 5 (a)) consistently exhibit clear cluster structures that concentrate on the target object and effectively distinguish it from the background.
Cross-Paradigm Visualization Comparison. Fig. 6 presents a comparative visualization analysis of different paradigms on the Mini-ImageNet classification task. For clustering-based models (CoC, ClusterFormer, FEC and CLUENet), we directly visualize semantic clusters at each stage. For FEC and our proposed CLUENet, we additionally show the final-stage clustering maps to illustrate their global semantic representations. For convolution-based models (e.g., ResNet), we use Grad-CAM (ref29) to highlight class-discriminative regions. For attention-based models (e.g., PVTv2), we visualize the attention response to the center key in the first attention head of the last block at each stage. We can see that ResNet primarily focuses on the head of the house finch, while PVTv2 attends to the torso. Both CoC and ClusterFormer produce fragmented clusters due to local window constraints, lacking coherent global semantics. In contrast, FEC and CLUENet consistently exhibit clear foreground-background separation across stages, indicating stronger interpretability. Moreover, CLUENet combines high accuracy, efficiency, and transparency, surpassing FEC in overall performance.
@@
Table 3: Ablation Studies on Core Components, Cluster Pooling, Query Design, and Model Variants
@@
We validates the effectiveness of model components on the Mini-ImageNet image classification task.
Key Components: As shown in Table 3(a), we conducted ablation studies on five core components of the CLUENet model to evaluate their individual contributions. Specifically, these components include: Feature Aggregation (FA), Temperature-Scaled Cosine Attention (TCosAttn), Gated Residual (Gate), Shared Dispatching (Shared), and Learnable Positional Embedding (PosEmb). Since both TCosAttn and Gated Residual are embedded within the global and soft feature aggregation module, removing FA also disables the other two. This leads to a significant drop in accuracy (Top-1: –1.78%, Top-3: –1.21%), highlighting the importance of global and soft feature aggregation in semantic integration. Replacing TCosAttn with standard attention reduces Top-1 and Top-3 accuracy by 0.59% and 0.11%, respectively, indicating its advantage in modeling semantic hierarchy and improving cluster-to-pixel alignment. Removing the Gated Residual and using fixed weights results in a smaller drop (–0.29%, –0.45%), showing that dynamic gating better adapts to local feature injection. Disabling shared dispatching causes accuracy to decrease by 0.43% (Top-1) and 0.78% (Top-3), along with increased parameters and computation. Finally, removing the learnable positional embedding leads to a 0.54% drop in Top-1 and 0.66% in Top-3 accuracy, suggesting its role in enhancing spatial awareness. Overall, all five components contribute meaningfully to performance and work together to improve the model’s representation capacity.
Cluster Pooling: Table 3(b) compares various cluster pooling designs. The similarity mapping layer proj_f in FEC shows performance nearly identical to an identity mapping, indicating model redundancy. Even when both proj_f and proj_v layers learn jointly (Fig. 4b), accuracy gains are negligible. Our improved cluster pooling module achieves a notable Top-1 accuracy gain (78.75% vs. 78.07%) while reducing computational cost, validating the design’s efficiency. Increasing proj_v depth from two to three layers yields no improvement, thus, the two-layer MLP is selected for balance.
Multi-Head Semantic Query: Table 3(c) shows that cluster center–guided queries significantly improve performance (78.75% vs. 77.72%), demonstrating enhanced semantic understanding. Queries based on initial average pooling lack the flexibility to capture rich semantic structures in natural images (78.75% vs. 78.02%), showing limitations compared to our cluster-guided queries. Multi-head queries outperform single-head (78.75% vs. 78.41%), confirming stronger semantic capture.
@@ Conclusions
We proposed CLUENet, a novel visual model that balances strong performance with inherent interpretability. By introducing a Global Feature Clustering Block with efficient attention and shared assignment across the global scope, we improve the model performance, computational efficiency, and enable enhanced visualization quality of the clustering paradigm. The improved cluster pooling overcomes gradient vanishing while maintaining excellent performance and visualization quality. Extensive experiments on CIFAR-100 and Mini-ImageNet show that CLUENet achieves superior accuracy compared to existing clustering-based and mainstream models, while also providing clear and intuitive semantic interpretability. Future work will focus on enhancing semantic information flow across stages to improve robust recognition in complex scenarios.
Acknowlegements
This work was supported by the National Natural Science Foundation of China under Project Nos. 62572480, 62201600, 62201604, 62506371, 62522604 and 62476258.

### v2

@@ Introduction
Results on CIFAR-100. As shown in Table 1, in the convolution paradigm, ShuffleNetv2 achieves the best Top-1 accuracy of 71.09%. For the attention paradigm, PVTv2 leads with 70.77%. Among clustering paradigm, CLUENet stands out with a Top-1 accuracy of 76.55%, surpassing existing clustering model CoC by 4.63% and outperforming PVTv2 and ShuffleNetv2 by 5.78% and 5.46%, respectively.
Results on Mini-ImageNet. Table 2 summarizes the classification results on the Mini-ImageNet validation set. Under similar parameter scales, CLUENet (base) achieves the best performance among clustering-based models, surpassing CoC (medium), FEC (large), and ClusterFormer (tiny) by 4.05%, 3.11%, and 8.45% in Top-1 accuracy, and by 2.18%, 1.92%, and 4.21% in Top-3 accuracy, respectively. Without using windowing mechanisms while adopting a relatively dense cluster count, CLUENet (base) achieves an inference speed of 679.05 img/s, which is comparable to that of FEC (550.69 img/s) and CoC (605.67 img/s), demonstrating its efficient architecture and strong feature representation.
Compared with the Attention-based models, CLUENet (micro) outperforms the attention-based PVTv2 (b0) by 3.41% in Top-1 accuracy and 2.31% in Top-3 accuracy while using 0.42M fewer parameters. It also surpasses EfficientFormer (s0) by 0.83% in Top-1 and 1.12% in Top-3 accuracy with a similar parameter budget. CLUENet (tiny) outperforms EfficientFormer (s1) by 1.54% in Top-1 and 0.97% in Top-3 accuracy. Furthermore, CLUENet (small) surpasses PVTv2 (b1) and EfficientFormer (s2) by 3.92% and 1.74% in Top-1, respectively.
Compared with the convolution-based models, CLUENet (micro) achieves slightly higher Top-1 and Top-3 accuracy than ShuffleNetv2 (x1.5) (+0.36%, +0.43%), and CLUENet (tiny) outperforms ShuffleNetv2 (x2.0) and ResNet18 by 0.88% and 3.56% in Top-1 accuracy, and by 0.36% and 1.41% in Top-3 accuracy, respectively. Moreover, CLUENet (small) surpasses both ResNet18 and ConvNeXtv2 (N) by 4.54% and 6.45% in Top-1 accuracy, respectively.
@@
Model Visualization. The proposed CLUENet offers intrinsic interpretability through its architecture. We visualize the cluster receptive fields of CLUENet (tiny) across all four stages on the Mini-ImageNet classification task. As shown in Fig. 5, the model captures multi-level semantic structures at different stages, reflecting a hierarchical feature extraction process similar to human visual perception. For clarity, only one semantic head is visualized per stage. Additionally, we visualize the receptive field of each position on the $7\times 7$ feature map from the fourth stage to highlight the semantic areas most influential for classification. Since directly displaying all 49 clusters can be overwhelming, we adopt the $K$-Means-based cluster merging method proposed by Chen et al. (ref19) to reduce the number of visualized clusters. To preserve the interpretability purity of the network, we visualize the fourth-stage $7\times 7$ feature map with increasing cluster numbers by adjusting the K-Means merge parameter.
(a) The semantic clustering.
(b) The impact of cluster numbers.
@@
Figure 6: Visualization results of different paradigm models. ResNet shows Grad-CAM activation maps at four stages. PVTv2 presents attention maps based on the center query token in the first attention head of the last block at each stage. CoC, ClusterFormer and FEC visualize one attention head from one block per stage. FEC and CLUENet additionally provide the final-stage semantic clustering maps used for classification.
As illustrated in Fig. 5(a), CLUENet exhibits a progressive semantic understanding. Early stages (1–3) focus on the object’s overall contour and location, separating it from the background to form clear object boundaries. The deeper stages (3–4) progressively shift attention to finer and more discriminative regions critical for classification. For example, in the first row of Fig. 5(a), the model initially treats the bird and branch as a whole, but by Stage 3, it distinguishes the branch as background and focuses on the bird’s head and torso in Stage 4. As shown in Fig. 5(b), when the number of clusters is further increased, the same semantic object is split into multiple semantic parts, capturing finer details of the object. Importantly, all visualized samples are correctly classified, and the final-stage feature maps (the last column of Fig. 5(a)) consistently exhibit clear cluster structures that concentrate on the target object and effectively distinguish it from the background.
Cross-Paradigm Visualization Comparison. Fig. 6 presents a comparative visualization analysis of different paradigms on the Mini-ImageNet classification task. For clustering-based models (CoC, ClusterFormer, FEC and CLUENet), we directly visualize the semantic clusters at each stage. For FEC and our proposed CLUENet, we additionally show the final-stage clustering maps to illustrate their global semantic representations. For convolution-based models (e.g., ResNet), we use Grad-CAM (ref29) to highlight class-discriminative regions. For attention-based models (e.g., PVTv2), we visualize the attention response to the center query in the first attention head of the last block at each stage. It is observed that ResNet converges on the head, while PVTv2 shifts focus from the torso to the head. Both CoC and ClusterFormer produce fragmented clusters due to local window constraints, thereby lacking coherent global semantics. In contrast, FEC and CLUENet consistently exhibit clear foreground-background separation across stages, indicating stronger interpretability. Moreover, CLUENet combines high accuracy, efficiency, and transparency, surpassing FEC in overall performance.
@@
Table 3: Ablation Studies of Core Components, Cluster Pooling, Query Design, and Model Variants
@@
We validate the effectiveness of model components on the Mini-ImageNet image classification task.
Key Components: As shown in Table 3(a), we conducted ablation studies on five core components of the model to evaluate their individual contributions. Specifically, these components include: Feature Aggregation (FA), Temperature-Scaled Cosine Attention (TCosAttn), Gated Residual (Gate), Shared Dispatching (Shared), and Learnable Positional Embedding (PosEmb). Since both TCosAttn and Gated Residual are embedded within the global and soft feature aggregation module, removing FA also disables the other two. Disabling FA leads to a significant drop in accuracy (Top-1: –1.78%, Top-3: –1.21%), highlighting the importance of global and soft feature aggregation in semantic integration. Replacing TCosAttn with standard attention reduces Top-1 and Top-3 accuracy by 0.59% and 0.11%, respectively, indicating its advantage in modeling semantic hierarchy and improving cluster-to-pixel alignment. Removing the Gated Residual and using fixed weights results in a smaller drop (–0.29%, –0.45%), showing that dynamic gating better adapts to local feature injection. Disabling shared dispatching causes accuracy to decrease by 0.43% (Top-1) and 0.78% (Top-3), along with increased parameters and computation. Finally, removing the learnable positional embedding leads to a 0.54% drop in Top-1 and 0.66% in Top-3 accuracy, suggesting its role in enhancing spatial awareness. Overall, all five components contribute meaningfully to performance and work together to improve the model’s representation capacity.
Cluster Pooling: Table 3(b) compares various cluster pooling designs. The similarity mapping layer $\operatorname{proj_{f}}$ in FEC shows performance nearly identical to an identity mapping, indicating model redundancy. Even when both $\operatorname{proj_{f}}$ and $\operatorname{proj_{v}}$ layers learn jointly (Fig. 4b), accuracy gains are negligible. Our improved cluster pooling module achieves a notable Top-1 accuracy gain (78.75% vs. 78.07%) while reducing computational cost, validating the design’s efficiency. Increasing $\operatorname{proj_{v}}$ depth from two to three layers yields no improvement, thus, the two-layer MLP is selected for balance.
Multi-Head Semantic Query: Table 3(c) shows that cluster center–guided queries significantly improve performance (78.75% vs. 77.72%), demonstrating enhanced semantic understanding. Queries based on initial average pooling lack the flexibility to capture rich semantic structures in natural images (78.75% vs. 78.02%), showing limitations compared to our cluster-guided queries. Multi-head queries outperform single-head queries (78.75% vs. 78.41%), confirming stronger semantic capture.
@@ Conclusions
We proposed CLUENet, a novel visual model that balances strong performance with inherent interpretability. By introducing a Global Feature Clustering Block with efficient attention and shared assignment across the global scope, we improve performance, computational efficiency, and visualization quality. The improved cluster pooling overcomes gradient vanishing while maintaining excellent performance and model interpretability. Extensive experiments on CIFAR-100 and Mini-ImageNet show that CLUENet achieves superior accuracy compared to existing clustering-based and mainstream models, while also providing clear and intuitive semantic interpretability. Future work will focus on enhancing semantic information flow across stages to improve robust recognition in complex scenarios.
Acknowledgements
This work was supported by the National Natural Science Foundation of China under Project 62572480, 62201600, 62201604, 62506371, 62522604 and 62476258.